American Products plans to lay off 10% of its employees in its aerospace division and 15% of its employees in its agricultural division. If altogether 12% of the 3000 employees in these two divisions will be laid off, then how many employees are in each division?
Answer by ankor@dixie-net.com(22740)

American Products plans to lay off 10% of its employees in its aerospace division and 15% of its employees in its agricultural division. If altogether 12% of the 3000 employees in these two divisions will be laid off, then how many employees are in each division?
:
Let x = no. aero em
Let y = no. ag em
:
x + y = 3000
x = 3000 - y
:
.10x + .15y = .12(3000)
.10x + .15y = 360
;
Substitute (3000-y) for x, solve for y:
.10(3000-y) + .15y = 360
:
300 - .10y + .15y = 360
.05y = 360 - 300
.05y = 60
y = 60/.05
y = 1200 ags get the ax
:
x = (3000-1200) = 1800 aeros join them
:
:
Check:
.10(1800) + .15(1200) =
180 + 180 = 360 which is 12% of 3000